OpenStudy (anonymous):

hey guys what is e^2ln3

OpenStudy (anonymous):

is it equal to 0

OpenStudy (anonymous):

maybe i should edit it : Find the exact value of e\[e^{2\ln 3}\]

OpenStudy (anonymous):

e is a constant

OpenStudy (sgadi):

\[e^{2\ln3}=e^{\ln3^2}=3^2=9\]

OpenStudy (anonymous):

bingo, thank you